OpenStudy (anonymous):

Solve. 1/2 = 5^x + 4

OpenStudy (anonymous):

x = negative 7 over 2 x = –6 x = 9 over 2 x = 2

OpenStudy (anonymous):

subtract 1/2 by 4 and divide by 5

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

that is raised to the x

sam (.sam.):

1/2 = 5^x + 4 Subtract 4 both sides 1/2 = 5^x + 4 -7/2=5^x ln(-7/2)=xln(5) x=ln(-7/2)/ln(5)